OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 a)−1 b)zero c)x can be any number as there are infinitely many solutions to this system d)There is no x value as there is no solution to this system

OpenStudy (anonymous):

Have you tried it or not??

OpenStudy (anonymous):

Yes i got a but i'm not sure if that is correct

mathslover (mathslover):

wht r u getting ?

OpenStudy (anonymous):

A..

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

I am not getting this..

OpenStudy (anonymous):

Check your solution once again.. Can you show your work??

mathslover (mathslover):

Should i show my work ?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Wait @mathslover

mathslover (mathslover):

k .. i will let u do it on ur own

mathslover (mathslover):

firstly find the value of x in the first equation 2x-3y=3

mathslover (mathslover):

in terms of y. . wht r u getting ?

OpenStudy (anonymous):

When i isolated y i for 2/3x

OpenStudy (anonymous):

got*

mathslover (mathslover):

no dont isolate y .. find the value of x ..

OpenStudy (anonymous):

Use elimination @mathslover it will be easy...

mathslover (mathslover):

\[\large{2x-3y=3}\] \[\large{2x=3+3y}\] divide both sides by 2 wht r u getting now ?

OpenStudy (anonymous):

x= 3/2 + 3/2y

OpenStudy (anonymous):

I tried using elimination @waterineyes and i got -1 :o

mathslover (mathslover):

good now put this value of x in 2nd equation that is : \[\large{5x-4y=4}\] \[\large{5(\frac{3+3y}{2})-4y=4}\] wht r u getting now ? yes u r right y = -1

mathslover (mathslover):

either do this in this way also ..if u want

OpenStudy (anonymous):

Okay... Firstly multiply the first equation by 4 and tell me what you got.

OpenStudy (anonymous):

-4(2x-3y=3) 3(5x-4y=4) it will give something like -8x=-12 & 15x=12 after combining both it will give 7x=0 x=0

OpenStudy (anonymous):

@mathslover question is demanding for x value and not y.. \(x \ne -1\)

OpenStudy (anonymous):

Multiply first equation by 4 and second by 3: \[8x - 12y = 12\] \(15x - 12y = 12\) \(-7x = 0\) \(x = 0\)

OpenStudy (anonymous):

Oh okay so x=0? Based on elimination i got 8x-12y = 12 for 1st equation -15x+ 12 = -12 for 2nd equation so 7x = 0 so x is 0?

OpenStudy (anonymous):

Thank you all for your help :)

OpenStudy (anonymous):

Yes you are right @helpmexoxo

mathslover (mathslover):

@waterineyes I had showed the relation between y and x and had calculated the value of y then one can easily calculate for x.............. @helpmexoxo Nice to hear that you got the correct way of answering the question............ :)

OpenStudy (anonymous):

Then why you said yes you are right y = -1 ?? @mathslover

mathslover (mathslover):

y=-1 and then x=(3+3y)/2=3-3/2=0/2=0

OpenStudy (anonymous):

Anyways thanks to you too for helping @helpmexoxo ..

mathslover (mathslover):

got it now ?

mathslover (mathslover):

Not thanks to me ... one who did correct deserves thanks